# Inside you are many wolves: Using cognitive models to interpret value trade-offs in LLMs

### v1

@@ Abstract
Navigating everyday social situations often requires juggling conflicting goals, such as conveying a harsh truth, maintaining trust, all while still being mindful of another person’s feelings. These value trade-offs are an integral part of human decision-making and language use, however, current tools for interpreting such dynamic and multi-faceted notions of values in LLMs are limited. In cognitive science, so-called “cognitive models” provide formal accounts of these trade-offs in humans, by modeling the weighting of a speaker’s competing utility functions in choosing an action or utterance. In this work, we use a leading cognitive model of polite speech to interpret the extent to which LLMs represent human-like trade-offs. We apply this lens to systematically evaluate value trade-offs in two encompassing model settings: degrees of reasoning “effort” in frontier black-box models, and RL post-training dynamics of open-source models. Our results highlight patterns of higher informational utility than social utility in reasoning models, and in open-source models shown to be stronger in mathematical reasoning. Our findings from LLMs’ training dynamics suggest large shifts in utility values early on in training with persistent effects of the choice of base model and pretraining data, compared to feedback dataset or alignment method. We show that our method is responsive to diverse aspects of the rapidly evolving LLM landscape, with insights for forming hypotheses about other high-level behaviors, shaping training regimes for reasoning models, and better controlling trade-offs between values during model training.
@@ 1 Introduction
People regularly contend with the goals and values of others. But people also regularly contend with competing goals and values within themselves. This inner goal conflict has been studied formally in philosophy, economics, AI, and cognitive science [e.g. 58, 1, 73, 15]. It is also a familiar aspect of how people intuitively describe their inner lives [^1] To give an example: in an often-repeated story, a person is told that inside them there is a battle between two wolves, one representing anger and malice, the other representing hope and kindness. When the person asks which wolf will win, they are told ‘the one you feed’.. This inner goal competition is present in major decisions, and it also suffuses everyday social communication. Even the simple act of telling your friend their cake is a disaster can require balancing your value of conveying the truth, with your value for your friend’s mental state and emotions. Such competing inner goals drive how people choose what to communicate, and the understanding of this competition is necessary for decoding what people mean from what they say and do.
A large body of work in cognitive science has formalized pragmatic communication in humans as a family of recursive probabilistic generative models, known as Rational Speech Acts (RSA) models [19, 23]. This class of cognitive models includes a pragmatic speaker that chooses what to say by balancing a mixture of goals (including being informative, but also various other affective, relational, and persuasive goals), and a pragmatic listener that interprets the speaker’s utterances and actions by taking into account such possible goals [e.g. 44, 45, 6, 7, 80].
@@
Ideally, conversational agents—including large language models (LLMs)—should exhibit similar sensitivity to human-like value trade-offs in communication. Yet, as decades of work has emphasized, endowing artificial agents with such nuanced social reasoning remains a foundational challenge [97, 14, 57]. While the current paradigm of value alignment has made considerable progress [40], there is reason to question whether guiding the output of models towards singular attributes like “helpfulness” or “truthfulness” can equip them with the representations needed to capture such trade-offs [52, 18].
@@
Here, we expand upon the growing toolkit of intepretability methods aimed at understanding the multifaceted nature of values in alignment [e.g. 88, 99, 51, 35], with cognitively interpretable models that are informed by value trade-offs made by the humans that interact with these systems. Our approach is grounded in an Inverse Reinforcement Learning (IRL) view of RLHF: namely, reverse-engineering the objectives that are implicit in human-provided behavior [89, 43]. We combine this view with theoretical connections to Theory-of-Mind inference in humans [38, 37], and suggest using cognitive models of pragmatic inference in humans to formalize evaluations of LLMs’ learned reward functions.
Figure 1: Paradigm overview. (1) We collected LLMs’ responses in a polite speech task, and fit a well-established probabilistic generative model of the behavior from Yoon et al., [92] to these data. (2) We report the results of the following inferred parameters of this model for two suites of LLMs: $\phi$, which describes the first-order speaker’s weighting of informational and social utilities, and $\omega$, which describes the second-order speaker’s weighting of informational, social, and presentational utilities. (3) A schematic illustration of the cognitive model of polite speech.
@@ 1.1 Contributions
We focus on the domain of polite language, as formalized by Yoon et al., [92] for two reasons: First, this domain naturally captures trade-offs between the kinds of opposing utilities that are central to the alignment problem in LLMs: how to convey true and useful information, while providing responses that are agreeable to human users. The importance of this particular set of value trade-offs has also recently been underscored by increasing concerns about sycophantic behavior in popular LLMs that prioritize pleasing a user over maintaining truthfulness [53, 64, 56]. Second, the communicative nature of the experimental stimuli used in Yoon et al., [92], more closely approximates the features of real-world LLM use cases compared to similar reference game tasks [50]. We apply this tool to a variety of closed and open-source large language models, and demonstrate the relevance of a structured probabilistic model of cognitive processes as a distinctive method for model interpretation.
To systematically study LLMs’ value trade-offs, we carefully designed two model suites for assessing a range of relevant model characteristics (see Table 1). Our closed-source model suite consists of three families of frontier models and three degrees of reasoning “effort” (from no reasoning to “medium effort”). Our open-source model suite is designed to disentangle the roles of model family, feedback dataset, and alignment method in the RL post-training process. We infer the parameters of the cognitive model over training checkpoints for a total of 8 unique configurations of these aspects. [^2] Code and data are available at https://github.com/skmur/many-wolves.
@@
Our results in these domains highlight patterns of higher informational than social utility in the reasoning variants of closed-source models, and in open-source models that are purported to be stronger in mathematical reasoning. Further, models’ training dynamics over the alignment process reveal that the largest shifts in utility values happen within the first quarter of training. Still, it appears that the choice of base model and pretraining data may have an outsized impact on the resulting weighting of utilities compared to the choice of feedback dataset or alignment method. Taken together, our findings suggest that this method is responsive to diverse aspects of the rapidly evolving LLM landscape: our tool provides opportunities for forming fine-grained hypotheses about other high-level behavioral concepts, understanding the extent of training needed to achieve particular values, and shaping recipes for higher-order reasoning and alignment capabilities.
2 Background
2.1 Value alignment in LLMs
A substantial body of work on aligning large language models (LLMs) has focused on optimizing models to reflect human preferences. Reinforcement learning-based methods—such as Reinforcement Learning from Human Feedback (RLHF) [78, 65, 4] and Reinforcement Learning from AI Feedback (RLAIF) [5]—as well as offline preference optimization techniques like Direct Preference Optimization (DPO) and variants [72, 16, 31, 68], have become standard components of the LLM alignment pipeline. These methods are widely believed to underlie many of the human-like behaviors exhibited by current models [40]. While off-policy methods and the use of static datasets are more efficient and easy to implement, prior work has shown that online methods are superior for preference learning [81, 82, 90]. However, prior work has also shown that the resulting models after preference fine-tuning generally show a lack of linguistic and conceptual diversity, which suggests a difficulty in maintaining multiplicity [47, 36, 66, 67, 61, 59, 87].
Recently, reinforcement learning-based finetuning has become popular for improving mathematical reasoning and coding abilities in models, where rewards are verifiable as opposed to coming from a learned reward model [93, 49, 34, 28, 76, 83]. Such ‘reasoning models’ exhibit certain characteristics such as having longer and more expressive chains of thought [86]. However, it is unclear what model behavior is elicited— even unintentionally— as a result of optimizing the verifiable rewards in these constricted domains; for instance, DeepSeek R1 underwent an additional stage of preference finetuning for safety alignment [28]. In spite of this, subsequent work has indicated that these reasoning models exhibit safety degradation [98, 33, 42].
2.2 Inverse RL for understanding agent behavior
A key limitation of the current RL*F paradigm is the opacity of the underlying learned reward function, which poses challenges for the safety and interpretability of the resulting model. Engineering reward functions that accurately describe real-world domains is nontrivial [2, 48]. One avenue for addressing this challenge has emerged from Inverse Reinforcement Learning (IRL), which seeks to infer a reward function from demonstrations provided by experts. Like RLHF, IRL aims to learn desired behavior from human input, but does so from expert demonstrations rather than preference feedback [45]. This connection suggests that IRL provides a useful conceptual and methodological lens for understanding and analyzing RLHF systems. In particular, IRL offers tools for interpreting and probing learned reward models by reconstructing the objectives implicit in human-provided behavior [89, 43].
Simultaneously, theory of mind and pragmatic inference in humans can also be thought of as a form of IRL in everyday social cognition. People regularly infer the goals and intentions of others from observed actions and utterances, providing a theoretical bridge between RLHF and the cognitive models that formalize these inferences in humans [37, 38]. These cognitive models offer another potential ground truth or benchmark for evaluating the robustness of learned reward functions under varying cognitive assumptions.
Model Family Model Path Reasoning Effort
Closed Models Anthropic claude-3-5-sonnet-20241022 None
[…]
Medium
[…]
Medium
[…]
Medium
Model Feedback Dataset Alignment Method
Open Models Qwen (Qwen2.5-7B-Instruct) HuggingFaceH4/ultrafeedback_binarized DPO
PPO
[…]
PPO
[…]
PPO
[…]
PPO
2.3 Using cognitive models to understand LLM behavior
Prior work has explored using the mathematical formalism of cogntive models to interpret the behavior of LLMs in a variety of settings [e.g. 74]. In the domain of pragmatic communication [27], prior work has characterized the goodness-of-fit of LLM behavior to different aspects of the Rational Speech Acts model [19]. Carenini et al., [8] considers the LLM as a listener in this model, while Jian and N, [41] explore methods for constructing the space of alternative utterances and meaning functions needed for RSA-based evaluations of LLMs. Of particular relevance to the alignment setting is [60], which proposes that RLHF post-training equips LLMs with a Theory-of-Mind-like abilities to anticipate a listener’s interpretation in its calculation of an output distribution.
The present work most closely relates to that of Liu et al., [54], which uses a cognitive model of trade-offs between honesty and helpfulness to evaluate LLMs in a signaling bandits experimental paradigm [79]. We extend the ideas in this work across a few dimensions. Firstly, we consider a related model of polite speech [92], which models opposing trade-offs between informational, social, and presentational goals in the task of giving feedback to someone in socially sensitive situations. While still a toy domain, this ungrounded, open-ended experimental paradigm better approximates the features and utilities of the alignment problem in LLMs. In addition to interpreting the behavior of black-box models, we also conduct a systematic analysis of these value trade-offs as a function of different base models, feedback datasets, and alignment methods in the RL post-training alignment process. Zhao and Hawkins, [95] also use this cognitive model of polite speech to investigate linguistic strategies in humans and LLMs in recent work, complementing our alignment-focused model analyses.
2.4 Reinforcement learning post-training dynamics
Several studies have examined how model behavior changes during reinforcement learning-based post-training, with the goal of understanding the specific contributions of RL relative to factors such as dataset composition and choice of base model. These studies have primarily focused on the setting of RL-based post-training for enhancing the mathematical reasoning and coding abilities of models  [96, 94] using verifiable rewards [49]. Of particular relevance is Gandhi et al., [20], which uses controlled behavioral evaluations to show that different base models exhibit varying degrees of reasoning behaviors—such as verification and backtracking—following RL post-training. The present work similarly leverages cognitive models to analyze the dynamics of RL post-training, but focuses on how LLMs implicitly learn more complex reward functions in an open-ended language domain where binary notions of “correctness” are not well-defined.
In the value alignment setting, prior work has analyzed the training dynamics of RLHF [21] and DPO [71], highlighting the issue of reward overoptimization—where proxy reward scores continue to improve while actual response quality stagnates or declines. Similarly, Chen et al., [10] identify limitations in both RLHF and DPO, showing that metrics such as ranking accuracy and win rate correlate positively only when the trained model remains close to the reference model.
@@ 3 Cognitive model
In this work, we consider the computational cognitive framework of polite speech production from Yoon et al., [92], an extended model in the Rational Speech Act framework [23]. This choice of domain is particularly relevant to value alignment, as it is pervasive, well-studied, and involves a fundamental trade-off between informational utility and social utility.
@@
The informativeness and the expected social value of an utterance $u$ are both a function of how the literal listener $L_{0}$ interprets utterances $P_{L_{0}}(s|u)$, which is grounded out on the literal semantics $[\![u]\!](s)$ with a prior over the states $s$ likely to be communicated, i.e. $P_{L_{0}}(s|u)\propto[\![u]\!](s)\cdot P(s)$. For simplicity, the mapping from true state $s$ (i.e. the speaker’s actual assessment of the listener’s creation, specified in terms of the number of stars they would give it; see Section 5.1) to its perceived social value, $V(s)$, is assumed to be an identity function.
Yoon et al., [92] fit the parameters of this model to interpret the structure underlying complex pragmatic behaviors in humans, and in this work, we do the same to understand LLMs’ behavior (see Section 5.2 and Section B.2 for details). The particular parameters of interest are $\phi$ and $\boldsymbol{\omega}$. As illustrated above, the mixture parameter $\phi$ captures the trade-off between informational and social utilities that the second-order pragmatic speaker $S_{2}$ wishes to project towards a lower-order pragmatic listener $L_{1}$. $\phi=1$ indicates high projected informational utility, while $\phi=0$ indicates high projected social utility. The trade-off ratios $\boldsymbol{\omega}$ captures how the second-order pragmatic speaker balances informational, social, and presentational goals.
@@ 4 Language model evaluation suites
We design two model suites for evaluation that cover a range of characteristics that are thought to have implications for LLMs’ ability to capture human-like value trade-offs (see Table 1).
Closed-source model suite
The objective of our closed-source model evaluations is two-fold. First, we aim to more rigorously interpret claims about the behavioral tendencies of widely-used black-box models. Second, we seek to understand how reasoning-optimized variants–models trained via extended RLHF to produce longer, more structured chains of thought [86], often for coding and math–might be adapting LLM behaviors in everyday contexts where value alignment is critical [cf. 98, 33, 42]. To these ends, we evaluate three degrees of reasoning in Anthropic, Google, and OpenAI’s models: a) models that do not explicitly use any additional chain-of-thought reasoning (Claude-Sonnet-3.7 [3], Gemini-Flash-2.0 [24], and ChatGPT-4o [62]), and b) the low and medium effort reasoning modes of their reasoning counterparts (Claude-Sonnet-3.7 [3], Gemini-2.5-Flash [25], o4-mini [63]). For Gemini and o4, these effort levels can be specified directly by the parameters low and medium, but for Claude-Sonnet-3.7, which instead uses a specific token count, we map these values to 1k tokens and 8k tokens, respectively, following the values indicated in the Gemini API documentation.
Open-source model suite
To understand which factors most influence model behavior after preference fine-tuning, we systematically evaluate the effects of base model family, preference dataset, and alignment algorithm on the resulting value trade-offs. Each of these elements —the pretraining distribution of the base model, the structure of the feedback dataset, and the choice of learning algorithm— has been shown to shape downstream behavior. For instance, Qwen models [91] are known to be pretrained on large amounts of synthetic data, especially in mathematical domains, in contrast to Llama [26]. Similarly, the Anthropic HH-RLHF dataset [4] emphasizes harmlessness and helpfulness, whereas UltraFeedback [12] contains more diverse instruction-following preferences. Recent work also suggests that the choice of alignment method can also impact outcomes, with PPO shown to induce less reward overoptimization compared to DPO [71]. The influence of each of these factors on learned value trade-offs remains unclear, motivating our controlled study of model checkpoints from combinations of the aforementioned models, datasets, and alignment methods. For each configuration (8 total), we initialize from an instruction-tuned model, perform one epoch of supervised fine-tuning (SFT) on the ‘chosen’ responses, and follow with one epoch of preference optimization using either DPO or PPO (implemented using OpenRLHF [32]) with ArmoRM [85] as the reward model. We evaluate each model’s behavior across evenly spaced checkpoints throughout the preference fine-tuning stage to trace the evolution of alignment and value trade-offs (see Section B.1 for full hyperparameter details).

### v3

# Using cognitive models to reveal value trade-offs in language models

@@ Abstract
Value trade-offs are an integral part of human decision-making and language use, however, current tools for interpreting such dynamic and multi-faceted notions of values in LLMs are limited. In cognitive science, so-called “cognitive models” provide formal accounts of such trade-offs in humans, by modeling the weighting of a speaker’s competing utility functions in choosing an action or utterance. Here we use a leading cognitive model of polite speech to systematically evaluate value trade-offs in two encompassing model settings: degrees of reasoning “effort” in frontier black-box models, and RL post-training dynamics of open-source models. Our results highlight patterns of higher informational utility than social utility in reasoning models’ default behavior, and demonstrate that these patterns shift in predictable ways when models are prompted to prioritize certain goals over others. Our findings from LLMs’ training dynamics suggest large shifts in utility values early on in training with persistent effects of the choice of base model and pretraining data, compared to feedback dataset or alignment method. Our framework offers a flexible tool for probing value trade-offs across diverse model types, providing insights for generating hypotheses about other social behaviors such as sycophancy and for shaping training regimes that better control trade-offs between values during model development.
@@ 1 Introduction
People regularly contend with the goals and values of others. But people also regularly contend with competing goals and values within themselves. This inner goal conflict has been studied formally in philosophy, economics, AI, and cognitive science (e.g. Minsky, 1986; Ainslie, 2001; Schelling et al., 1984; Dennett, 1991), is present in major decisions, and suffuses everyday social communication. Even the simple act of telling your friend that their cake is a disaster can require balancing your value of conveying the truth, with your value for your friend’s feelings. Such competing inner goals drive how people choose what to communicate, and the understanding of this competition is necessary for decoding what people mean from what they say and do.
@@
Ideally, conversational agents—including large language models (LLMs)—should exhibit similar sensitivity to human-like value trade-offs in communication. Yet, as decades of work has emphasized, endowing artificial agents with such nuanced social reasoning remains a foundational challenge (Zhi-Xuan et al., 2024; Dennett, 1987; McCarthy, 1979). While the current paradigm of value alignment has made considerable progress (Ji et al., 2024), there is reason to question whether guiding the output of models towards singular attributes like “helpfulness” or “truthfulness” can equip them with the representations needed to capture such trade-offs (Lindström et al., 2024; Fish et al., 2025).
A large body of work in cognitive science has formalized pragmatic communication in humans as a family of recursive probabilistic generative models, known as Rational Speech Acts (RSA) models (Frank & Goodman, 2012; Goodman & Frank, 2016).
Figure 1: We use cognitive models that are designed to explain the structure of human behavior (top) to interpret how low-level training decisions impact LLMs’ representations of human-like value trade-offs (bottom).
This class of cognitive models includes a pragmatic speaker that chooses what to say by balancing a mixture of goals (including being informative, but also various other affective, relational, and persuasive goals), and a pragmatic listener that interprets the speaker’s utterances and actions by taking into account such possible goals (e.g. Kao et al., 2014; Kaufmann et al., 2024; Barnett et al., 2022; Carcassi & Franke, 2023; Sumers et al., 2024).
@@
Here, we expand upon the growing toolkit of intepretability methods (e.g. Wollschläger et al., 2025; Zou et al., 2025; Lindsey et al., 2025; Jain et al., 2024) with cognitive models that are designed to explain the structure of human-generated behavioral data. We use this to investigate the extent to which pluralistic value structures emerge through alignment. Since LLMs are trained on precisely such data, we posit that cognitive models offer a valuable ground truth or benchmark for evaluating the robustness of learned reward functions as a result of lower-level modeling decisions. Our approach is grounded in an Inverse Reinforcement Learning (IRL) view of RLHF: namely, reverse-engineering the objectives that are implicit in human-provided behavior (Wulfmeier et al., 2024; Joselowitz et al., 2025). We combine this view with theoretical connections to Theory-of-Mind inference in humans (Jara-Ettinger et al., 2016; Jara-Ettinger, 2019), and suggest using cognitive models of pragmatic inference in humans to formalize evaluations of LLMs’ learned reward functions (see Appendix A for detailed background).
@@ 1.1 Contributions
We focus on doing so in the domain of polite language, as formalized by Yoon et al. (2020) for two reasons: First, this domain naturally captures trade-offs between the kinds of opposing utilities that are central to the alignment problem in LLMs: how to convey true and useful information, while providing responses that are agreeable to human users. The importance of this particular set of value trade-offs has also recently been underscored by increasing concerns about sycophantic behavior in popular LLMs that prioritize pleasing a user over maintaining truthfulness (Liu et al., 2025; OpenAI, 2025; Marks et al., 2025). Second, the communicative nature of the experimental stimuli used in Yoon et al. (2020), more closely approximates the features of real-world LLM use cases compared to similar reference game tasks (Lewis, 1969).
We apply this tool to a variety of closed and open-source large language models (see Appendix Table 1), and demonstrate the relevance of a structured probabilistic model of cognitive processes as a distinctive method for model interpretation. Our closed-source model suite consists of three families of frontier models across three values of reasoning budget (none, low, and medium), with analyses of the effects of prompt-based manipulations that simulate different “goals” a speaker can have (to be informative, social, or both). Our open-source model suite is designed to disentangle the roles of model family, feedback dataset, and alignment method in the RL post-training process. We infer the parameters of the cognitive model over training checkpoints for a total of 8 unique configurations of these aspects.
@@
Our results highlight patterns of higher informational utility than social utility in reasoning models’ default behavior, and demonstrate that these patterns shift in predictable ways when models are prompted to prioritize certain goals over others. Further, models’ training dynamics over the alignment process reveal that the largest shifts in utility values happen within the first quarter of training. Still, it appears that the choice of base model and pretraining data may have an outsized impact on the resulting weighting of utilities compared to the choice of feedback dataset or alignment method. Taken together, our findings suggest that this method is responsive to diverse aspects of the rapidly evolving LLM landscape: our tool provides opportunities for forming fine-grained hypotheses about other high-level behavioral concepts, understanding the extent of training needed to achieve particular values, and shaping recipes for higher-order reasoning and alignment capabilities.
@@ 2 Cognitive model
In this work, we consider the computational cognitive framework of polite speech production from Yoon et al. (2020), an extended model in the Rational Speech Act framework (Goodman & Frank, 2016). This choice of domain is particularly relevant to value alignment, as it is pervasive, well-studied, and involves a fundamental trade-off between informational utility and social utility.
@@
The informativeness and the expected social value of an utterance $u$ are both a function of how the literal listener $L_{0}$ interprets utterances $P_{L_{0}}(s|u)$, which is grounded out on the literal semantics [[u]](s)[\![u]\!](s) with a prior over the states $s$ likely to be communicated, i.e. $P_{L_{0}}(s|u)\propto[\![u]\!](s)\cdot P(s)$. For simplicity, the mapping from true state $s$ (i.e. the speaker’s actual assessment of the listener’s creation, specified in terms of the number of stars they would give it; see Section 4.1) to its perceived social value, $V(s)$, is assumed to be an identity function.
Yoon et al. (2020) fit the parameters of this model to interpret the structure underlying complex pragmatic behaviors in humans, and in this work, we do the same to understand LLMs’ behavior (see Section 4.2 and Section C.2 for details). The particular parameters of interest are $\phi$ and $\boldsymbol{\omega}$. As illustrated above, the mixture parameter $\phi$ captures the trade-off between informational and social utilities that the second-order pragmatic speaker $S_{2}$ wishes to project towards a lower-order pragmatic listener $L_{1}$. $\phi=1$ indicates high projected informational utility, while $\phi=0$ indicates high projected social utility. The trade-off ratios $\boldsymbol{\omega}$ captures how the second-order pragmatic speaker balances informational, social, and presentational goals.
Figure 2: Paradigm overview. (1) We collected LLMs’ responses in a polite speech task, and fit a well-established probabilistic generative model of the behavior from Yoon et al. (2020) to these data. (2) We report the results of the following inferred parameters of this model for two suites of LLMs: $\phi$, which describes the first-order speaker’s weighting of informational and social utilities, and $\omega$, which describes the second-order speaker’s weighting of informational, social, and presentational utilities. (3) A schematic illustration of the cognitive model of polite speech.
@@ 3 Language model evaluation suites
We design two model suites for evaluation that cover a range of characteristics that are thought to have implications for LLMs’ ability to capture human-like value trade-offs (see Appendix Table 1). In the closed-source setting, we aim to understand the behavioral tendencies of widely-used black-box models and how their reasoning-optimized variants might be adapting LLM behaviors in everyday contexts where value alignment is critical (cf. Zhou et al., 2025; Huang et al., 2025; Jiang et al., 2025). In the open-source setting, we seek to understand which factors influence model behavior after preference fine-tuning by systematically evaluating the effects of base model family, preference dataset, and alignment algorithm on the resulting value trade-offs.
Closed-source model suite We evaluate three degrees of reasoning in Anthropic, Google, and OpenAI’s models: a) models that do not explicitly use any additional chain-of-thought reasoning (Claude-Sonnet-3.7, Gemini-Flash-2.0, and ChatGPT-4o), and b) the low and medium effort reasoning modes of their reasoning counterparts (Claude-Sonnet-3.7, Gemini-2.5-Flash, o4-mini). For the Gemini and o4 models, these effort levels can be specified directly by the parameters “low” and “medium”, but for Claude-Sonnet-3.7, which instead uses a specific token count, we map these values to 1k tokens and 8k tokens, respectively, following the values indicated in the Gemini API documentation.
Open-source model suite We consider all unique configurations of two 7B parameter base models (Qwen2.5-Instruct (Yang et al., 2024) and Llama-3.1-Instruct (Grattafiori et al., 2024), two feedback datasets (UltraFeedback and Anthropic HH-RLHF  (Bai et al., 2022a), and two learning algorithms (DPO (Rafailov et al., 2024) and PPO). These elements have been shown to have characteristics that shape downstream behavior, however the influence of each of these factors on learned value trade-offs remains unclear. For each configuration (8 total), we initialize from an instruction-tuned model, perform one epoch of supervised fine-tuning (SFT) on the ‘chosen’ responses, and follow with one epoch of preference optimization using either DPO or PPO (implemented using OpenRLHF (Hu et al., 2024)) with ArmoRM (Wang et al., 2024) as the reward model. We evaluate each model’s behavior across evenly spaced checkpoints throughout the preference fine-tuning stage to trace the evolution of alignment and value trade-offs (see Section C.1 for full hyperparameter details).